How do I write a quadratic equation with 9 as the only solution?
Answer by Edwin McCravy(20054)



     x = 9;     x = 9

 x - 9 = 0; x - 9 = 0

   (x - 9)(x - 9) = 0

x2 - 9x - 9x + 81 = 0 

    x2 - 18x + 81 = 0

Edwin